# PatchTrack: A Comprehensive Analysis of ChatGPT’s Influence on Pull Request Outcomes

### v1

@@ 3.2.4 Step 4: Pull request classification
The classification of pull requests in Step 3 follows a straightforward approach. For each class, we aggregate its occurrences within a pull request. If the count of PA is greater than zero, the entire pull request is classified as PA. If there are no instances of PA but there are instances of PN, the pull request is classified as PN. If no ChatGPT code snippets were identified in step 3, both PA and PN counts for this pull request will be zero, leading to an overall classification of NE. The analysis provided by PatchTrack shows how the 645 ChatGPT-generated code snippets were compared with 3,486 pull request patches, resulting in precise classifications and valuable insights into the interaction between developers and AI-generated patches.
3.2.5 Step 5: Compute percentage of ChatGPT patch applied
Since we are performing an exact match in clone detection, we employed the Jaccard containment metric, defined as:
Where and represent the token streams from ChatGPT and GitHub patches, respectively. This metric calculates the proportion of common code by dividing the number of tokens at the intersection of and by the total number of tokens in the ChatGPT suggested patch. This metric was adopted from the original Redebug paper [53]. The result shows the percentage of ChatGPT’s code integrated into the pull request by the developer.
3.2.6 Step 6: Qualitative analysis
To address RQ2, RQ3, and RQ4, we employed a rigorous qualitative analysis methodology. This involved summarizing ChatGPT developer conversations and GitHub pull requests using the Framework Method [54], followed by thematic extraction using the card sorting technique [55]. This process was carried out collaboratively by 2 Ph.D. students and 1 Professor from the same lab.
Step 6.1: Summarization using the Framework Method:
We collected relevant ChatGPT–developer conversations and GitHub pull requests categorized as PN and NE, organized into a spreadsheet. In the pull requests, we specifically focused on the sections where the ChatGPT link was shared, examining the code and developer interactions a few steps before and after the link to ensure relevance to the ChatGPT conversation, as the pull requests might contain unrelated content.
Each team member familiarized themselves with the data, and we developed an initial coding framework based on our research questions, focusing on elements such as the nature of the advice given by ChatGPT, the type of challenge encountered by the developer, and the reasons for non-integration of patches or absence of patch suggestions.
Through iterative discussions, we refined this framework and systematically applied it to the data, creating summaries for each interaction. These summaries captured key information such as the developer’s problem, the advice provided by ChatGPT, how the developer applied or did not apply this advice, and the outcomes of the interaction.
To enhance the quality of our summaries, we leveraged ChatGPT version 4.0, providing it with organized ChatGPT developer conversations and the corresponding pull requests for each instance. We compared ChatGPT’s summaries with our manual ones, finding ChatGPT’s results often superior in English context and clarity. However, in instances where ChatGPT produced hallucinations, we resorted to our manually constructed summaries. For inter-coder reliability, each team member independently summarized a subset of cases, and we reached a consensus on the final summaries.
Step 6.2: Card Sorting for Theme Extraction:
Using the summaries created in the Framework Method, we prepared detailed cards for each interaction. Each team member individually sorted the summary cards into categories based on perceived similarities and emerging themes, grouping the cards into thematic clusters. After individual sorting, we conducted a group discussion to compare and consolidate the categories, refining and agreeing on the final set of themes. We then performed a thematic analysis on the agreed categories to identify and define the overarching themes, reviewing them for coherence, refining them to ensure that they accurately represented the data, and naming them appropriately.
4 Results & Discussion - RQ1
@@
How are ChatGPT-generated patches (PA, PN, and NE) distributed and utilized in merged pull requests under self-admitted ChatGPT usage?
This research question quantifies the distribution of pull requests classified as PA, PN, and NE within our 285-case dataset.
To support this classification, we evaluated the tokenization strategy used to compare ChatGPT-suggested code with pull request patches. Since ChatGPT-generated snippets are often brief, selecting an appropriate n-gram size is critical to capturing exact matches while minimizing false positives. We conducted a comparative analysis across n-gram values , and , validating results with both automated metrics and manual inspection.
4.1 Token Matching Methodology
Our evaluation showed that consistently achieved the best performance, accuracy (97.5%), precision (96.5%), recall (96.7%), and F1-score (96.6%), particularly for detecting short and exact matches. Larger values reduced recall and did not capture concise code fragments. Manual validation (Cohen’s Kappa = 0.85) further confirmed that provided the most reliable alignment between the suggestions generated by ChatGPT and the patches applied by the developer.
We also examined how token size influenced patch classification outcomes. Smaller n-grams (n = 1) identified more patch applied (PA) cases and fewer patch not applied (PN) cases, whereas larger n-grams were more sensitive to formatting and syntactic variations, leading to under-detection. The complete results are provided in the appendix A.
Misclassification Analysis
We identified and manually analyzed 7 misclassified cases (4 patch not applied PN and 3 patch applied PA), to better understand PatchTrack’s limitations under the setting. Two key factors contributed:
1.
Inaccessible Patches via GitHub API: GitHub does not expose file diffs exceeding 400 lines via its API. As a result, large patches were omitted during extraction. In three cases, ChatGPT-generated code was present in these omitted sections, causing false PN classifications (e.g., PR #2232 [56]).
2.
Single-Line Matches: While the tool correctly flags one-line matches, this sometimes led to misclassifying a PN case as PA. These borderline cases highlight a trade-off between sensitivity and semantic depth.
Overall, while PatchTrack effectively identifies directly applied patches, it does not capture structural or semantic transformations. Because it relies on token-level matching rather than higher-order semantic analysis (e.g., Type-4 clone detection), cases with significant refactoring.
4.2 Patch Classification
We analyze 645 ChatGPT-generated code snippets and 3,486 patches in 285 pull requests. To systematically classify these snippets, we employed PatchTrack, which compares each ChatGPT-generated snippet with all patches in its corresponding pull request. This approach enabled us to categorize the snippets into predefined integration groups. Through a series of experiments, we found that using an n-gram value of achieved the highest accuracy, precision, recall, and F1 score compared to other sizes of n-grams. This choice was guided by the concise nature of the ChatGPT-generated patches, where exact token matching improves detection accuracy while minimizing the risk of overlooking minor but significant similarities (see Table 2–appendix A and Section 4.1 for a detailed analysis). Figure 5 summarizes the classification results, illustrating the distribution of pull requests in PA, PN, and NE.
4.2.1 Percentage of ChatGPT Patches Applied
@@
To measure the integration of suggested ChatGPT patches into pull requests, we used Jaccard’s containment ratio 1, which quantifies the overlap between suggested patches and the final code integrated into pull requests. Figure 6 illustrates the distribution of integration percentages for the classification PA. The median integration rate is 25%, indicating that, on average, a quarter of the ChatGPT-suggested code is adopted in merged pull requests.
@@
To better understand this variation, we selected four pull requests representing different levels of integration, with one pull request chosen from each quartile: (i) low (0–25%], where only minor elements of the suggested patch were retained, often requiring substantial modifications or alternative implementations; (ii) lower-median (25–50%], where developers incorporated portions of the patch but refined it significantly; (iii) upper-median (50–75%], where a larger portion of the patch was adopted with some modifications; and (iv) high (75–100%], where the AI-generated patch was largely accepted with minimal changes. Each quartile follows the standard convention where the lower bound is excluded, and the upper bound is included. These cases were selected through stratified random sampling to ensure balanced representation across integration levels. By analyzing these strata, we illustrate real-world scenarios of ChatGPT-assisted patch integration, shedding light on why some patches are fully applied, while others are selectively merged or heavily revised.
@@
Example 1: Low Patch Integration (0–25%] In pull request #185[57] of the pokt-network/poktroll repository, the developer asked ChatGPT111Example 1 link to ChatGPT conversation – Accessed: 2025-02-17 if Go had a built-in function to remove leading spaces from multiline strings while maintaining code readability. ChatGPT suggested an approach using strings.Split, strings.TrimLeft, and strings.Join to process each line and remove leading spaces. The developer then requested a refined solution that encapsulated this logic into a reusable helper function, trimString, to improve readability. ChatGPT provided a full Go program, but only its helper function was integrated. Of the suggested 31 LOC, just 7 LOC (22.6%) were adopted. The minimal adoption reflects the developer’s selective integration of only the core functionality needed, as the existing project structure did not require a complete rewrite.
Example 2: Lower-median Patch Integration (25% - 50%] pull request #7123 [58] in the Mudlet/Mudlet repository addressed a crash that occurred when double-clicking on a word in TTextEdit. The issue stemmed from an out-of-bounds access in a function responsible for expanding text selection, leading to an assertion failure in a QString. After encountering a stack trace pointing to the faulty function, the developer consulted ChatGPT222Example 2 link to ChatGPT conversation – Accessed: 2025-02-17, which identified the issue and suggested a structured fix: (i) Implement boundary checks for yind, (ii) Break loops when indices become invalid, and (iii) Refactor selection logic to improve robustness. The initial ChatGPT-generated patch (39 LOC) effectively prevented the crash by halting execution before an invalid index was accessed. However, only 18 LOC (46.2%) were ultimately integrated after a reviewer suggested negating conditions to reduce indentation and improve readability. The extent of the modifications suggests that while the ChatGPT recommendations provided a strong starting point, the developers refined the patch to better align with the project constraints and coding practices. This case illustrates how AI-assisted debugging can be useful but requires human oversight and adjustments to ensure code quality and maintainability.
@@
Example 3: Upper-median Patch Integration (50% - 75%] pull request #279 [59] in the nylas/nylas-python repository introduced support for a new “free-busy” API endpoint in the Nylas SDK, enabling users to check calendar availability for specific email addresses. A key challenge in this implementation was designing a suitable data model to represent the API response, which could include both availability data and error messages. To address this, the developer consulted ChatGPT333Example 3 link to ChatGPT conversation – Accessed: 2025-02-17 for guidance on structuring Python dataclasses for the JSON response. ChatGPT suggested: (i) defining separate data models for distinct response components, (ii) utilizing Python’s dataclass module for maintainability, and (iii) employing Union types to handle multiple object types within a single list. The proposed patch consisted of 25 lines of code (LOC), of which 15 LOC (60%) were integrated. The primary modification involved renaming the error class to freeBusyError for improved clarity, based on reviewer feedback. This case highlights how AI-generated code can closely align with project requirements, yet still requires refinements guided by domain expertise to enhance readability and maintainability.
@@
Example 4: High Patch Integration (75–100%] In pull request #2059 [60] of the alshedivat/al-folio repository, a developer extended a Jekyll plugin to support fetching posts from a manually specified list of URLs in addition to an RSS feed. Initially, the plugin retrieved posts only via RSS, but the developer aimed to enhance its functionality by incorporating direct URL input. To achieve this, they consulted ChatGPT444Example 4 link to ChatGPT conversation – Accessed: 2025-02-17 to modify the Ruby-based Jekyll plugin, focusing on three key aspects: (i) metadata extraction, (ii) correct handling of published dates, and (iii) improved logging readability. ChatGPT contributed by refactoring the code to introduce modular functions, ensuring robust date handling with Time objects, improving log clarity through puts, and enhancing maintainability by extracting date parsing into a dedicated parse_published_date function. The patch consisted of 19 lines of code (LOC), all of which were integrated as-is (100%), as it aligned perfectly with the project’s needs, requiring no modifications. The developer’s enthusiasm, expressed at the end of their interaction with ChatGPT (“Awesome! I have a feeling we’ll write a lot of code together from now on ;)”), underscores how ChatGPT’s response was fully applicable without requiring further adjustments.
4.3 Discussion and Implications - RQ1
Our findings indicate that while ChatGPT-generated patches provide value, full adoption remains rare, with a median integration rate of 25%. Developers often treat AI-generated suggestions as a starting point rather than a final implementation, selectively integrating elements that align with project-specific constraints. The third quartile (38%) further reinforces that even in cases with higher integration, significant modifications are frequently required. Integration patterns reveal different adoption behaviors: In low integration cases (0-25%), developers retained only minor portions of AI-generated code due to inefficiencies, redundancy, or misalignment with project conventions. Moderate integration cases (25–75%) show that developers adopted substantial sections but refined them to fit architectural guidelines, while high integration cases (75–100%) occurred when ChatGPT-generated patches closely matched project needs, requiring little or no modification.

### v2

@@ 3.2.4 Step 4: Pull request classification
The classification of pull requests in Step 3 follows a straightforward approach. For each class, we aggregate its occurrences within a pull request. If the count of PA is greater than zero, the entire pull request is classified as PA. If there are no instances of PA but there are instances of PN, the pull request is classified as PN. If no ChatGPT code snippets were identified in Step 3, both PA and PN counts for this pull request will be zero, leading to an overall classification of NE. Pull requests that were closed without merging are classified as CL, regardless of whether they contain ChatGPT-suggested patches, allowing us to analyze closure-specific factors in RQ4. The analysis provided by PatchTrack shows how the 645 ChatGPT-generated code snippets were compared with 3,486 pull request patches, resulting in precise classifications and valuable insights into the interaction between developers and AI-generated patches. While PatchTrack classifies each individual hunk, our analysis aggregates the results at the pull request level. This is because patch usage decisions (application, revision, rejection, or closure) are typically discussed and resolved within the PR context rather than at the level of individual snippets.
3.2.5 Step 5: Qualitative Analysis
To address RQ1–RQ4, we conducted a qualitative analysis of pull requests pre-processed and classified by PatchTrack. For RQs, we analyzed cases labeled PA, PN, NE, and CL, which represent scenarios where ChatGPT suggestions were adopted, modified, conceptually reused, or ultimately rejected. For each case, we examined the linked ChatGPT conversation, the surrounding GitHub discussion (up to three developer messages before and after the shared link), and relevant code diffs. This interaction window was selected based on pilot testing, which showed it typically captured sufficient context without including unrelated discussion. Code changes were also reviewed to assess alignment with AI suggestions.
We applied the Framework Method [52], proceeding through structured summarization, codebook development, and theme refinement. One author and a research assistant independently reviewed overlapping subsets of PR cases and generated structured summaries using a consistent template. Each summary captured the developer’s problem, the ChatGPT recommendation, the developer’s response, the resulting code outcome, and any reviewer feedback. To improve consistency and reduce fatigue for longer threads, we tested GPT-4-generated summary drafts. These were always reviewed and revised by a human analyst. In cases where the model introduced hallucinations or omitted relevant context, the human-authored version was prioritized. All summaries were then reviewed and finalized in discussion with a senior researcher to ensure agreement and interpretive alignment across cases.
All summaries were collaboratively coded and grouped into higher-level themes through card sorting, continuing the three-person team from the summarization phase (one author, a research assistant, and a senior researcher). Initial coding and grouping were performed by the first author and assistant, with discrepancies discussed and resolved in consensus meetings involving the second author and senior researcher. These discussions helped clarify the relationships between surface behaviors and underlying motivations, ensuring interpretive consistency. While we did not compute Cohen’s Kappa due to the qualitative nature of the task, agreement was verified through iterative discussion and convergence on shared meanings. The final thematic structure was applied across the full dataset.111The research assistant contributed only to manual analysis and theme sorting. They were not involved in study design, interpretation, or writing and therefore do not meet authorship criteria. This follows ACM’s authorship policy: https://www.acm.org/publications/policies/new-acm-policy-on-authorship
While our coding was primarily inductive, we consulted prior studies on pull request rejection and review dynamics [18, 15, 17] to inform our general understanding. These provided background perspective but did not directly guide our codes, as they do not account for the unique characteristics of AI-assisted contributions. Because AI-generated patches introduce distinct risks and social dynamics, such as hallucinated output, prompt ambiguity, developer over-reliance on AI, and uncertainty in trust attribution, we extended our codebook by open coding 20 representative cases to surface patterns grounded in the data. Examples of structured summaries and the code-to-theme mapping are included in our replication package [42] to support transparency and reproducibility.
4 Results and Discussion
This section presents the key findings of our study. We begin by characterizing the repositories and pull requests that form the basis of our dataset. This provides essential context for interpreting how ChatGPT-generated patches are used in real-world development workflows. We then answer each of our research questions in turn, starting with RQ1, which quantifies the integration of ChatGPT-generated code across pull requests.
4.1 Repository Characteristics
The 338 pull requests in our dataset come from 255 unique GitHub repositories that vary widely in maturity, popularity, and technical focus. Figure 3 summarizes core repository attributes. Popularity indicators are long-tailed: stars range from 0 to 81,000 (median: 58) and forks from 0 to 36,655 (median: 31). Commit histories are likewise highly skewed, with counts peaking above 1.2 million (median: 946). The language panel further indicates coverage across diverse implementation ecosystems. These attributes indicate that our dataset includes both niche and high-profile projects. Prior studies have used similar repository-level metrics to capture ecosystem diversity and developer participation [53, 54, 55].
In terms of language diversity, the dataset covers 123 distinct programming languages. The most frequently observed languages include Shell (85.5%), JavaScript (82.0%), HTML (62.0%), TypeScript (50.0%), Python (54.1%), and CSS (54.0%), reflecting a strong representation of web development and scripting tasks. Infrastructure and configuration files are also well represented, such as Dockerfile (49.0%) and Makefile (33.0%). Compiled and system-level languages such as Java (19.0%), C (18.4%), and C++ (14.5%) are present but less dominant.
While this linguistic variety highlights the breadth of our dataset, its distribution diverges from GitHub’s overall language trends. According to GitHub’s 2024 Octoverse report [56], Python has surpassed JavaScript as the most used language, with Java, C, and C++ also ranking among the top. In contrast, our dataset overrepresents Shell, JavaScript, and other front-end and scripting technologies. This suggests that developers are more likely to use ChatGPT in tasks that involve scripting, web development, and configuration management, domains that favor automation and well-defined coding patterns. By contrast, areas that demand deeper domain knowledge or complex architectural decisions (e.g., backend systems, data science) may be less conducive to AI-assisted contributions.
These findings suggest that current ChatGPT usage in pull request workflows is concentrated in specific development contexts. Broader conclusions about generative AI adoption will require datasets with greater alignment to GitHub’s full language ecosystem. Future work could address this by stratifying repositories according to GitHub language distributions and exploring AI-supported development practices across a wider range of software domains.
4.2 RQ1: Classification of Patch Use in Pull Requests
@@
RQ1: How can PatchTrack classify the integration of ChatGPT-generated patches in pull requests, and what patterns emerge in how developers adopt these patches under self-admitted ChatGPT usage?
This research question investigates the distribution of ChatGPT-assisted pull requests across three integration categories: Patch Applied (PA), Patch Not Applied (PN), and No Patch Exists (NE). Using the final configuration of PatchTrack (), we analyzed 285 pull requests, resulting in the following classification:
Patch Applied (PA): 116 cases (40.7%)
Patch Not Applied (PN): 63 cases (22.1%)
No Patch Exists (NE): 106 cases (37.2%)
Figure 4 visualizes this distribution. While many developers adopted ChatGPT-generated patches, others rejected or heavily modified them, and in some cases, ChatGPT did not generate a patch at all. These outcomes reflect varying levels of trust, relevance, and integration effort, which we explore further in RQ2–RQ4.
4.2.1 Patch Classification
We analyzed 645 ChatGPT-generated code snippets and 3,486 patches across 285 pull requests. Each snippet was compared against all patches in its corresponding PR using PatchTrack, which employs token-level matching to determine integration. This process enabled classification into PA, PN, or NE, as described in Section 3.2.3.
As validated in our methodology (Section 3.2.3), using an n-gram size of yielded the highest accuracy, precision, recall, and F1 score, due to its sensitivity to short code fragments. This configuration was used for all reported results. Figure 4 summarizes the resulting distribution of PRs across the three integration categories.
Percentage of ChatGPT Patches Applied
To assess how much of ChatGPT’s suggested code is integrated into pull requests, we compute the percentage of ChatGPT patches applied. This metric helps evaluate the degree to which developers adopt AI-generated suggestions when resolving issues or adding features. By linking overlap percentage to classification labels, we can systematically evaluate developer adoption behavior and determine how self-admitted ChatGPT usage translates into actual code integration. For example, a high ratio indicates direct reuse of ChatGPT’s output, while a low ratio may suggest heavy modification or rejection. This directly supports our goal of classifying integration outcomes. We use Jaccard’s containment ratio 1 to quantify the overlap between tokens in the ChatGPT-suggested patch () and the final integrated code ():
@@ 4.2.1 Patch Classification
Here, and are token streams from ChatGPT and GitHub patches, respectively. This formulation, adopted from the Redebug study [57], measures the proportion of ChatGPT’s code that appears in the final merged pull request. Figure 5 illustrates the distribution of integration percentages for the classification PA. The median integration rate is 25%, indicating that, on average, a quarter of the ChatGPT-suggested code is adopted in merged pull requests.
@@
To better understand this variation, we selected four pull requests representing different levels of integration, with one pull request chosen from each quartile: (i) low (0–25%], where only minor elements of the suggested patch were retained, often requiring substantial modifications or alternative implementations; (ii) lower-median (25–50%], where developers incorporated portions of the patch but refined it significantly; (iii) upper-median (50–75%], where a larger portion of the patch was adopted with some modifications; and (iv) high (75–100%], where the AI-generated patch was largely accepted with minimal changes. Each quartile follows the standard convention where the lower bound is excluded and the upper bound is included. These cases were selected through stratified random sampling to ensure balanced representation across integration levels. By analyzing these strata, we illustrate real-world scenarios of ChatGPT-assisted patch integration, shedding light on why some patches are fully applied while others are selectively merged or heavily revised. These examples are illustrative and not intended for theory-building, so theoretical saturation was not a goal.
4.2.2 Qualitative Analysis – RQ1
To complement the quantitative analysis of patch adoption rates, we conducted a thematic analysis of the validated PA cases. Although 90 cases were initially sampled, four were later reclassified as PN and three from PN to PA, resulting in a final set of 89 PA cases (see Section 3.2.3 for details on misclassifications). Coding followed the structured framework described in our methodology. Agreement was reached through coder discussion and convergence on six themes. Concise theme definitions are provided in the replication package [42]; the main text focuses on their distribution and illustrative cases.
Table 3 summarizes how themes distribute across integration ranges. For brevity, we present one representative example per quartile below. Each instance is labeled PA-X (where X is the instance number). Full case summaries, coding assignments, pull request links, and ChatGPT–developer conversation links are available in the replication package [42].
| Theme | (0–25%] | (25–50%] | (50–75%] | (75–100%] | Total |
| Constraint Handling | 4 | 3 | 3 | 0 | 10 |
| Direct Implementation | 0 | 0 | 0 | 3 | 3 |
| Iterative Refinement | 10 | 3 | 7 | 6 | 26 |
| Knowledge Support | 4 | 1 | 5 | 3 | 12 |
| Selective Extraction | 8 | 9 | 0 | 1 | 18 |
| Structural Integration | 8 | 9 | 1 | 1 | 19 |
| Total | 34 | 25 | 16 | 14 | 89 |
@@ 4.2.2 Qualitative Analysis – RQ1
Low Integration (0–25%): This quartile (34 cases) is characterized by Structural Integration (8 cases) and Selective Extraction (8 cases) as the most distinctive behaviors for minimal carryover of AI code, even though Iterative Refinement appears slightly more often numerically (10 cases). We emphasize Structural Integration and Selective Extraction here because they most clearly explain why integration remained low: developers either reshaped ChatGPT’s output to fit project frameworks or extracted only fragments of functionality, discarding the rest. These patterns reflect how AI output typically required substantial reshaping before it could be meaningfully integrated in this range (counts from Table 3). Example 1 (PA-27)–Case ID: PA-27; Theme: Structural Integration illustrates this pattern: In pull request #185 [58] of the pokt-network/poktroll repository, the developer asked ChatGPT222Example 1 link to ChatGPT conversation – Accessed: 2025-02-17 if Go had a built-in function to remove leading spaces from multiline strings while maintaining code readability. ChatGPT suggested an approach using strings.Split, strings.TrimLeft, and strings.Join to process each line and remove leading spaces. The developer then requested a refined solution that encapsulated this logic into a reusable helper function, trimString, to improve readability. ChatGPT provided a full Go program, but only its helper function was integrated. Of the suggested 31 LOC, just 7 LOC (22.6%) were adopted. The minimal adoption reflects the developer’s selective integration of only the core functionality needed, as the existing project structure did not require a complete rewrite.
Lower-median Integration (25–50%): This quartile (25 cases) is dominated by Selective Extraction (9 cases) and Structural Integration (9 cases), with Constraint Handling and Iterative Refinement also present but less frequent (counts from Table 3). Developers in this range often incorporated substantial portions of ChatGPT’s suggestions, but not wholesale: they selectively retained useful fragments or adapted the AI output to align with coding conventions. Integration percentages therefore remained below 50%, reflecting the tendency to treat ChatGPT’s patch as scaffolding rather than final code. Example PA-36; Theme: Selective Extraction illustrates this pattern: In pull request #7123 [59] of the Mudlet/Mudlet repository, a crash occurred when double-clicking a word in TTextEdit, traced to an out-of-bounds access in text-selection logic. After reviewing the stack trace, the developer consulted ChatGPT333Example 2 link to ChatGPT conversation – Accessed: 2025-02-17, which proposed a structured patch: (i) add boundary checks for yind, (ii) break loops when indices become invalid, and (iii) refactor selection logic for robustness. The 39-LOC patch addressed the crash, but only 18 LOC (46.2%) were integrated. A reviewer suggested simplifying the control flow by negating conditions to reduce indentation, prompting further edits. The adopted code retained the essential boundary checks while discarding or refactoring the rest, showing how developers extracted only the useful fragments of ChatGPT’s logic while reshaping them to improve readability and project alignment.
@@
Upper-median Integration (50–75%): This quartile (16 cases) is characterized by Iterative Refinement (7 cases) as the most common theme, followed by Constraint Handling (3 cases) and Knowledge Support (5 cases) (counts from Table 3). Developers in this range integrated a majority of ChatGPT’s suggestions, but usually with targeted refinements rather than direct adoption. Reviewer input often drove these refinements, such as clarifying class names, reorganizing logic, or improving readability. Integration percentages in this quartile therefore reflect substantial reuse of ChatGPT code, tempered by human oversight to align with project conventions. Example PA-73; Theme: Iterative Refinement illustrates this pattern: Pull request #279 [60] in the nylas/nylas-python repository introduced support for a new “free-busy” API endpoint in the Nylas SDK, enabling users to check calendar availability for specific email addresses. A key challenge was designing a suitable data model to represent the API response, which could include both availability data and error messages. To address this, the developer consulted ChatGPT444Example 3 link to ChatGPT conversation – Accessed: 2025-02-17 for guidance on structuring Python dataclasses. ChatGPT suggested: (i) defining separate data models for distinct response components, (ii) using Python’s dataclass module for maintainability, and (iii) employing Union types to handle multiple object types within a list. The proposed patch contained 25 LOC, of which 15 LOC (60%) were integrated. The main adjustment involved renaming the error class to freeBusyError for clarity, following reviewer feedback. This case shows how AI-generated code can align closely with project needs while still requiring human-led refinements to enhance readability and maintainability.
@@
High Integration (75–100%): This quartile (14 cases) is dominated by Direct Implementation (3 cases) alongside instances of Iterative Refinement (6 cases) and Knowledge Support (3 cases) (counts from Table 3). Developers in this range adopted the ChatGPT output with little modification, often because AI-generated code already aligned with project requirements. When changes occurred, they were minor and stylistic rather than structural. These cases highlight scenarios in which ChatGPT produced production-ready code that required minimal oversight before integration. Example PA-82; Theme: Direct Implementation illustrates this pattern: Pull request #2059 [61] of the alshedivat/al-folio repository extended a Jekyll plugin to support fetching posts from a manually specified list of URLs in addition to an RSS feed. Initially, the plugin retrieved posts only via RSS, but the developer aimed to enhance functionality by incorporating direct URL input. To achieve this, they consulted ChatGPT555Example 4 link to ChatGPT conversation – Accessed: 2025-02-17, focusing on three key aspects: (i) metadata extraction, (ii) correct handling of published dates, and (iii) improved logging readability. ChatGPT refactored the Ruby plugin to introduce modular functions, ensured robust date handling with Time objects, improved log clarity through puts, and extracted date parsing into a dedicated parse_published_date function. The patch consisted of 19 LOC, all of which were integrated as-is (100%), since the output aligned perfectly with project needs. The developer’s enthusiasm, expressed at the end of the interaction (“Awesome! I have a feeling we’ll write a lot of code together from now on ;)”), underscores how ChatGPT’s response was fully applicable without requiring further adjustments.
4.2.3 Discussion and Implications – RQ1
Our findings show that while ChatGPT-generated patches provide value, full adoption remains uncommon, with a median integration rate of 25%. Developers most often treated ChatGPT’s output as a starting point rather than a final implementation, selectively integrating elements that aligned with project-specific needs. Thematic analysis of 89 validated PA cases revealed four dominant integration patterns across quartiles: Selective Extraction, Structural Integration, Iterative Refinement, and Direct Implementation. These patterns explain not only how much of ChatGPT’s code was reused but also why integration remained partial or complete.
In low-integration cases (0–25%), developers retained only minor portions of ChatGPT’s suggestions, typically extracting fragments or reshaping code to fit project frameworks. Moderate integration cases (25–75%) revealed substantial reuse, but usually with targeted refinements, such as simplifying control flow, renaming classes, or aligning with architectural constraints, often informed by reviewer feedback. High-integration cases (75–100%) occurred when ChatGPT’s output already conformed to project conventions, enabling near-direct adoption with minimal oversight. These variations highlight that ChatGPT accelerates exploration but rarely produces production-ready patches without human adaptation.